OpenStudy (anonymous):

Write the exponential equation in logarithmic form. x^0 = 1, (x is different than 0)

OpenStudy (anonymous):

\[\ln 1=0\]

OpenStudy (ash2326):

We have \[x^0=1\] Take log both sides \[\log_{10} x^0= \log_{10} 1\] We know \[\log_a b^c=c \times\log_a b\] so \[\log_{10} x^0= \log_{10} 1\] We get \[0 log_{10} x=\log_{10} 1\] so \[0=0\] This is a trivial equation any value of x>0 will satisfy this